OpenStudy (anonymous):

What is the axis of symmetry for the graph of the following function? y = -2(x + 4)^2 + 7 The axis of symmetry is x =

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (amistre64):

what makes the x part go to zero?

OpenStudy (amistre64):

x+4 = 0 gives us the axis of symmetry

OpenStudy (anonymous):

so x = 0?

OpenStudy (amistre64):

lets chk ... does 0+4 = 0 ??? no

OpenStudy (anonymous):

no but 4 + -4=0

OpenStudy (amistre64):

x + 4 = 0 when x=-4, therefore the axis of symmetry is x=-4

OpenStudy (amistre64):

\[\large y = -2~\underbrace{(x + 4)}_{\text{helps define axis of symmetry}}~^2 + 7\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do i need to do

OpenStudy (amistre64):

you might want to start at the top and read thru it ... the answer has already been given

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Can you help me with question?